OpenStudy (anonymous):

Use distributive property to find the product. (3x-2)(5x+7)

OpenStudy (anonymous):

foil this; 3x*5x=15x^2 3x*7=21x -2*5x=-10x -2*7=-14 ==15x^2+1x-14

OpenStudy (campbell_st):

so its 15x^2 + 21x -10x -14 collect like terms 15x^2 + 11x -14

OpenStudy (anonymous):

mine should be 11x lol